OpenStudy (anonymous):

use crammers rule to solve for x, y and z in: x - 5y + 3z = 15 2x + 4y +7z = 34 x - y = 15 + 2z

OpenStudy (anonymous):

Why not just row reduce?

OpenStudy (anonymous):

i want a step-by-step solution. preparing for an exam tomorrow and kind of a bit saturated now. help please.

OpenStudy (anonymous):

A step by step solution to calculating a determinant?

OpenStudy (anonymous):

to solve a 3x3 matrix.

OpenStudy (anonymous):

what are the answers to x, y and z?

OpenStudy (anonymous):

You do know that in all cases row reduction is much faster. If your teacher isn't asking you to use a specific method you should row reduce.

OpenStudy (anonymous):

he mentioned crammers rule. how do u row-reduce wt crammers rule?

OpenStudy (anonymous):

It's explained step by step in the link I gave u...

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Gaussian_elimination

OpenStudy (anonymous):

He want's Cramer's rule...

OpenStudy (anonymous):

pls help confirm Im correct so far: D= -74, Dx = 248, Dy= -20 AND Dz = -16

OpenStudy (anonymous):

yes, he says to use crammers rule to solve.

OpenStudy (anonymous):

Fine

OpenStudy (anonymous):

I agree it's a waste of time and energy...

OpenStudy (anonymous):

Learn crammers rule but tell your teacher that its inefficient and he should accept the answer by using row reduction.

OpenStudy (anonymous):

my solution?

OpenStudy (anonymous):

Now just divide each one by D...

OpenStudy (anonymous):

he gave the question when he was teachin us matrices in class.

OpenStudy (anonymous):

was my solution correct?

OpenStudy (anonymous):

I assume so, I am not going to check simple numerical calculations......

OpenStudy (anonymous):

anyways, i appreciate your time. but how do you row-reduce?

OpenStudy (anonymous):

What did you get for x, y, z?

OpenStudy (anonymous):

U can enter matrices into Wolfram like this to check determinants http://www.wolframalpha.com/input/?i=Det+%28%281%2C-5%2C3%29%2C%282%2C4%2C7%29%2C%281%2C-1%2C2%29%29

OpenStudy (anonymous):

What did you get for x, y, z? It is up above...

OpenStudy (anonymous):

x=-3.35 y=0.27 z=0.22

OpenStudy (anonymous):

\[\left[ \begin {array}{cccc} 1&0&0&{\frac {581}{37}} \\0&1&0&{\frac {10}{37}}\\ 0&0&1& {\frac {8}{37}}\end {array} \right] \]

OpenStudy (anonymous):

Your solution is not correct.

OpenStudy (anonymous):

i guessed as much.

OpenStudy (anonymous):

this looks complicated.